OpenStudy (anonymous):

MEDAL AND FAN!!!!! Geraldo’s current salary is $40,000 per year. His annual pay raise is always a percent of his salary at the time. What would his salary be if he got four consecutive 4% increases? a. $46,794.34 c. $43,328.04 b. $52,697.23 d. $62,156.12

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

\[P(t)=M*(1+i)^t\] \[M=40,000\\i=4\%=0.04\\t=4\] \[P(t)=40,000*(1+0.04)^t\] \[P(4)=40,000*(1.04)^4\] \[\boxed{P(4)=46,794.34}\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You're welcome ;)